# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

### v3

@@ 5 Experiments
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$.
  • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights.
@@
  • Gc-train-to-Gc-infer: Unlike node-level tasks, we can infer on $G^{\prime}$ (Gc-infer) itself for graph-level tasks after training on $G^{\prime}$. Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
@@
[…]
Tasks Datasets
Node Regression
(Train=30%, Val=20%, Test=50%)
Chameleon [20]
Crocodile [20]
Squirrel [20]
Node Classification
(Train=20*c, Val=30*c, Test=Rest)
Cora [21]
Citeseer [22]
Pubmed [23]
Coauthor Physics [24]
DBLP [25]
OGBN-Products [26]
Graph Regression
(Train=50%, Val=25%, Test=25%)
QM9 [27]
ZINC (Subset) [28]
Graph Classification
(Train=50%, Val=25%, Test=25%)
PROTEINS [29]
AIDS [29]
Table 2: The table shows the datasets used and the splits in each dataset. ‘c’ represents the number of classes for the dataset. We applied the public split “fixed” from [30] for Cora, Citeseer, and PubMed. For the “random” split in node classification, we keep 2020 nodes per class for training, $30$ nodes for validation, and the rest as test nodes. [31] grouped the 1919 properties of each molecule in QM9 dataset into four broad categories. We predict one property from each broad category: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
@@ 6 Result and Analysis
Methods Models
Reduction
Ratio (r)
Dataset
@@
6.1 Model Performance
Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$
@@ 6.1 Model Performance
Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE) with coarsening ratios $0.1$. Results using different coarsening ratios are in Table 12. As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, an increase in the coarsening ratio decreases performance.
Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 13. Results show that the accuracy is comparable to the baselines. Table 5 shows results on the OGBN-Products dataset. Result on training on the full dataset from the OGBN-Products leaderboard [Link]. For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing. For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. More results at Table 14. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high. It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Table 7 shows graph classification results. Training and inferring $G^{\prime}$ outperforms the baselines.
Methods Models
Reduction
Ratio (r)
Dataset
@@
Table 6: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
@@
Table 7: Results on Graph Classification (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
@@ 6.2 Inference Time and Memory
Figure 3: GPU memory consumption (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference.
In Table 8(a), we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges. We see up to $\mathbf{100\times}$. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach can give an inference on the whole graph. Table 11 in the Appendix shows the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the baseline.
Table 8(b) compares the inference time of graph classification and graph regression tasks. We predict for randomly selected $1000$ graphs from the test split. The table shows how our method is comparable and sometimes faster than the baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
@@ Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) [8]
@@ Appendix C More on Time and Space complexity
Figure 4: The plot shows the feasibility region (green) corresponding to the coarsening ratio and Number of nodes (In order of Millions) in the graph for a fixed feature dimension of $d=200$.
In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities. Figure 4 shows that those conditions can be satisfied for any dataset with nodes in order of millions.
For node-level tasks, we show that the inference phase will take less time for single node prediction. To show that the time difference between the time taken by the classical approach and ours will increase as the number of nodes increases, we can formulate the time difference as follows:
$\mathcal{T}(n)=\underset{\textup{Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}-\underset{\textup{ FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (3)
In the above equation, we use the notations already defined in Section 4.3. We can also assume that $\max_{i}^{k}n_{i}=\frac{n}{\alpha}$, where $\alpha>1$ is some constant. Then we can rewrite the Equation 3 as:
$\mathcal{T}(n)=n^{2}d+nd^{2}-(\frac{n}{\alpha}+\phi_{\max})^{2}d+(\frac{n}{\alpha}+\phi_{\max})d^{2}$
We can observe that the above equation is strictly increasing for $n>0$. Hence the time difference also increases as the number of nodes in a dataset increases.
Table 9: Upper Bound of $\phi_{\textup{max}}$ values for different datasets and coarsening ratios.
Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
PROTEINS $6.999$ $0.333$ 1.000-1.000 1.571-1.571
AIDS $322.000$ $82.000$ $34.000$ $13.429$
QM9 $79.000$ $19.000$ $7.000$ $1.857$
ZINC (subset) 11.000-11.000 4.333-4.333 3.000-3.000 2.429-2.429
@@ Appendix F Results
Table 11 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Also, in Table 12 we show how the FIT-GNN model performs on Node Regression datasets for different coarsening ratios with Cluster Nodes method under Gs-train-to-Gs-infer setup.
Dataset Appending Nodes FIT-GNN Baseline
$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
Extra Node $0.622$ $0.622$ $0.622$ $0.564$
Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
Extra Node $4.586$ $4.586$ $4.586$ $4.586$
Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
Extra Node $7.190$ $7.190$ $7.190$ $7.190$
Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
Extra Node $1.906$ $1.103$ $0.970$ $0.970$
Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
Extra Node $30.097$ $2.490$ $1.835$ $1.835$
Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
Extra Node $0.778$ $0.562$ $0.562$ $0.562$
DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
Extra Node $28.271$ $6.008$ $2.308$ $2.207$
Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
Extra Node $41.066$ $19.936$ $19.936$ $19.936$
OGBN-Products Cluster Node $28.399$ $2840.706$
Methods Models Reduction Ratio (r) Dataset
Chameleon Crocodile Squirel
Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
Table 12: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
@@
Table 13: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
Methods Models Reduction Ratio (r) Dataset
ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $\mathbf{0.524}$
0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
0.5 $0.620$ $\mathbf{0.693}$ $1.227$ $0.475$ $0.561$
0.7 $0.658$ - - - -
SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
0.3 $0.671$ - - - -
0.5 $0.695$ - - - -
0.7 $0.645$ - - - -
GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
0.3 $0.669$ - - - -
0.5 $0.631$ - - - -
0.7 $0.658$ - - - -
Table 14: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Table 15: The Tables show the comparison between Extra Nodes and Cluster Nodes methods for node classification and node regression task. We show the average of the top 10 accuracies (higher the better) of the node classification task on Cora dataset and the average of the best 10 losses (lesser the better) of the node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup. They demonstrate the superior performance of the Cluster Nodes method over Extra Nodes.
(a) Node Classification on Cora.
r Extra Nodes Cluster Nodes
$0.1$ $\mathbf{0.798\pm 0.008}$ $0.790\pm 0.004$
$0.3$ $0.798\pm 0.008$ $\mathbf{0.800\pm 0.003}$
$0.5$ $0.807\pm 0.004$ $\mathbf{0.829\pm 0.002}$
$0.7$ $0.797\pm 0.007$ $\mathbf{0.813\pm 0.002}$
(b) Node Regression on Chameleon.
r Extra Nodes Cluster Nodes
0.1 $\mathbf{0.504\pm 0.004}$ $0.506\pm 0.006$
0.3 $0.579\pm 0.006$ $\mathbf{0.489\pm 0.003}$
0.5 $0.585\pm 0.004$ $\mathbf{0.484\pm 0.004}$
0.7 $\mathbf{0.504\pm 0.006}$ $0.539\pm 0.009$
Table 16: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
@@
Table 17: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
Cora Chameleon
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$

### v4

@@ 5 Experiments
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$. Our aim was to examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
  • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights. We explore whether training solely on the $G^{\prime}$ is sufficient for inference on $\mathcal{G}_{s}$. This setting presents a computationally efficient option, allowing users with limited resources to obtain results close to the best reported.
@@
  • Gc-train-to-Gc-infer: Unlike node-level tasks, for graph-level tasks, inference can be done directly on $G^{\prime}$ because the label corresponds to the entire graph and not to individual nodes. Therefore, even with a reduced graph representation, meaningful inference is possible. Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
@@
Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: FIT-GNN.
Table 2: Datasets and corresponding splits used in our experiments. ‘c’ denotes the number of classes in node classification datasets. Public split “fixed” from 10.5555/3045390.3045396 is used for Cora, Citeseer, and PubMed. For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing. For QM9, we follow DBLP:journals/corr/GilmerSRVD17 and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
Task Dataset (Reference) Train Val Test
Node Regression Chameleon rozemberczki2019multiscale 30% 20% 50%
Crocodile rozemberczki2019multiscale 30% 20% 50%
Squirrel rozemberczki2019multiscale 30% 20% 50%
Node Classification Cora McCallum2000 $20\times c$ $30\times c$ Rest
Citeseer 10.1145/276675.276685 $20\times c$ $30\times c$ Rest
Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 $20\times c$ $30\times c$ Rest
Coauthor Physics DBLP:journals/corr/abs-1811-05868 $20\times c$ $30\times c$ Rest
DBLP 10.1145/1401890.1402008 $20\times c$ $30\times c$ Rest
OGBN-Products 10.5555/3495724.3497579 $20\times c$ $30\times c$ Rest
Graph Regression QM9 DBLP:journals/corr/WuRFGGPLP17 50% 25% 25%
ZINC (Subset) DBLP:journals/corr/Gomez-Bombarelli16 50% 25% 25%
Graph Classification PROTEINS Morris+2020 50% 25% 25%
AIDS Morris+2020 50% 25% 25%
@@ 6 Result and Analysis
6.1 Model Performance
Methods Models Reduction Ratio (r) Dataset
@@ 6.1 Model Performance
Node Regression: Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE). As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, an increase in the coarsening ratio decreases performance.
Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$. variation_neighborhoods coarsening algorithm is used. OOM means Out Of Memory.
@@
Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 10. Results show that the accuracy is comparable to the baselines. Figure 5 shows the comparison of performance between different experimental setups for different methods of appending nodes on the Cora dataset. From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task. For the ’None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance. We also observe in the ’None’ method that, with an increase in the coarsening ratio, the performance drops. This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss. For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases. This can be argued because of the change in the distribution of the coarsening graph and the set of subgraphs. When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance. While comparing the methods of appending nodes, we infer that the Cluster Node method is performing similarly compared to the Extra Nodes method, if not better. Table 5 shows results on the OGBN-Products dataset. Result on training on the full dataset from the OGBN-Products leaderboard [Link]. For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing. For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
Table 6: Results on Graph Regression with MAE as metric (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Methods Models Reduction Ratio (r) Dataset
@@
0.3 $0.669$ $0.723$ $0.761$ $0.509$ $0.854$
0.5 $0.631$ $0.725$ $0.848$ $0.737$ $1.058$
Graph Regression: Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large. It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Similarly, when inferred on a coarsened graph for each graph in the dataset, the performance degrades. This is because, when a graph is reduced to a smaller graph, the positional information is lost, which is crucial since in graphs representing different molecules, the position of a certain component in a molecule results in different properties of the molecule.
Table 7: Results on Graph Classification with accuracy as metric (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
@@
Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$. For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data. DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution. Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class. However, the graph size is kept similar to the graph size in the training set. These approaches also don’t cover the test set. Compared to these baselines, our approach reduces each graph into a set of subgraphs and a coarsened graph based on a reduction ratio. As a result, each graph in the training set and test set gets reduced. From our results, we see that FIT-GNN outperforms all the baselines.
Table 11 and 12 present the ablation study on various coarsening algorithms for all the aforementioned tasks. From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others. Therefore, we report results on all datasets using variation_neighborhoods coarsening algorithm.
@@ 6.2 Inference Time and Memory
Figure 3: GPU memory consumption in MegaBytes (MB) (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference. variation_neighborhoods coarsening algorithm is used.
In Table 8(a), we show the average time to predict for $1000$ queries (Prediction for Node $v_{i}$) in the Baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges. We see up to $\mathbf{100\times}$. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach can give an inference on the whole graph. Figure 3 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline. Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset. As the coarsening ratio increases, the time also increases, since the number of subgraphs increases with the coarsening ratio.
Table 8(b) compares the inference time of graph classification and graph regression tasks. We predict for randomly selected $1000$ graphs from the test split. The table shows how our method is comparable and sometimes faster than the Baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
@@ Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) (10.1145/3447548.3467256)
@@ Appendix C More on Time and Space complexity
In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities.
For node-level tasks, to obtain lower FIT-GNN inference time complexity as compared to classical GNN, we want the following as per asymptotic analysis,
  1. 1.
    Baseline vs. Single-Node Inference:
    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (4)
  2. 2.
    Baseline vs. Full-Graph Inference:
    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\sum_{i=1}^{k}\bigl[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}\bigr]}}$ (5)
Figure 4: The figure shows the feasibility of our inference methods for different node-level datasets against coarsening ratios. Two important inference setups for each dataset: single-node inference with the computational cost $\mathcal{O}\left(\max_{i}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})+n\right)$, and full-graph inference with computational cost $\mathcal{O}\left(\sum_{i=1}^{k}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})\right)$, are compared against the baseline (classical GNN) computational cost of $\mathcal{O}\left(n^{2}d+nd^{2}\right)$. The coarsening method used is variation_neighborhoods. Both axes are in logarithmic scale (base 10).
Figure 4 demonstrates the feasibility of our FIT-GNN method for different coarsening ratios with regards to full-graph inference time complexity. We empirically compute the LHS (baseline is classical GNN) and RHS of the Inequality 4 and 5for various coarsening ratios and plot the results for multiple datasets. The coarsening algorithm used is variation_neighborhoods.
We make the following inferences:
  1. 1.
    There is a clear dataset-dependent computational-cost trade-off across different coarsening ratios. Moreover, the minimum inference time for our approach occurs at different coarsening ratios for different datasets.
  2. 2.
    Under single-node inference setup, FIT-GNN demonstrates increasingly superior inference time performance as compared to classical GNN with increasing coarsening ratio. This is because the individual subgraph sizes decrease with increasing coarsening ratio.
  3. 3.
    We observe lower inference time complexity of Cluster_node method as compared to Extra_node method under both inference setups. This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.
@@ Appendix F Results
Figure 5 presents the ablation study conducted to compare the performance of FIT-GNN for different experimental setup using different methods for appending nodes in the subgraph. The experiments are conducted on the Cora dataset for different coarsening ratios.
Figure 5: The plot shows the ablation study conducted on the Cora dataset to determine which experimental setup performs better than the others. The plot also compares different methods of appending nodes to the subgraphs and how performance changes with varying coarsening ratios. variation_neighborhoods coarsening algorithm is used.
Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios. It is observed that the None method takes the least amount of time, which is obvious since the subgraphs are not appended with additional nodes. For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended. For Cluster Node, we observe the maximum time taken since not only are we looking at the neighboring nodes that are lost due to partition, but also we are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
Figure 6: The plot shows how the coarsening time varies for the Cora dataset for different coarsening ratios using different methods of appending nodes to the subgraphs. variation_neighborhoods coarsening algorithm is used.
Table 10 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$. To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 11 and 12. Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Note that this memory consumption is to store the graph and the weight matrix in the memory only.
@@
Cora Chameleon
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
Table 11: The table shows the FIT-GNN ablation study to compare various coarsening methods on Cora and Chameleon datasets, and the metrics for each dataset are accuracy (higher the better) and normalized MAE (lower the better), respectively.
@@
Table 12: The table shows the FIT-GNN ablation study to compare various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
Dataset Appending Nodes FIT-GNN Baseline
$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
Extra Node $0.622$ $0.622$ $0.622$ $0.564$
Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
Extra Node $4.586$ $4.586$ $4.586$ $4.586$
Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
Extra Node $7.190$ $7.190$ $7.190$ $7.190$
Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
Extra Node $1.906$ $1.103$ $0.970$ $0.970$
Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
Extra Node $30.097$ $2.490$ $1.835$ $1.835$
Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
Extra Node $0.778$ $0.562$ $0.562$ $0.562$
DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
Extra Node $28.271$ $6.008$ $2.308$ $2.207$
Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
Extra Node $41.066$ $19.936$ $19.936$ $19.936$
OGBN-Products Cluster Node $28.399$ $2840.706$